OpenStudy (anonymous):

I've done the work but I need help figuring out this last part. Can someone help?

OpenStudy (cggurumanjunath):

?

OpenStudy (anonymous):

If the parent function f(x) = x2 is horizontally compressed by a factor of 3, translated 8 units to the right, then translated 1 unit down, write the resulting function g(x) in vertex form. a. g(x) =1/3(x - 8)2 - 1 b. g(x) =1/3(x + 8)2 - 1 c. g(x) = 3(x - 8)2 - 1 d. g(x) = 3(x + 8)2 - 1